OpenStudy (anonymous):

Write an equation for the line (6,5)(-5,-3)

OpenStudy (aravindg):

Really?

OpenStudy (nottim):

Honestly, on my part, I haven't done math for a while.

OpenStudy (comm.dan):

Is this where you have to find the slope of a line? @ashleynicolee

OpenStudy (anonymous):

slope and y intercept

OpenStudy (comm.dan):

Do you remember the formula for finding the slope of a line? @ashleynicolee

OpenStudy (anonymous):

(8x+7)/11 y=mx+n where m is slope coming from (5-(-3))/(6-(-5)) to find n place points (6,5) to equation as (x,y) in y=8/11x+n then n=7/11